40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed Apr 3 16:23:26 2019, anchal, DailyProgress, PMC, Rehabilitation of PMCs 
    Reply  Thu Apr 4 15:33:34 2019, anchal, DailyProgress, PMC, Rehabilitation of PMCs 
Message ID: 2313     Entry time: Wed Apr 3 16:23:26 2019     Reply to this: 2314
Author: anchal 
Type: DailyProgress 
Category: PMC 
Subject: Rehabilitation of PMCs 

Since the last incident of power supply tripping, I tested all photodiodes and electronics in the lab one by one. During the process, I realized that we are using way too low power in our experiment. According to my calculations, the current RFPD are such that they will be shot noise limited only with more than 2.8 mW light falling on them.


North Path

  • The north path's PMC servo card was used with external mixers. I found that we were driving EOMs way too lightly with the modulation index of about 0.05 rad only.
  • The path from FPTest1 to the first summing junction had a MAX333A and a low pass filter of an arbitrary pole in it. We are already using an external in-line 19 MHz low pass filter after the mixer downconversion.
  • This path has something which has broken down as it is railing the rest of the circuit.
  • I switched to input point FPTest2 which directly goes to the inverting input of the summing junction. This however changed required LO phase delay by pi.
  • So I optimized the phase delay again and cut a RG-178 cable for it.
  • In the end, with 6.08 mW power falling on North PMC, I see PDH error signal of about 490 mVp-p with power going into EOM being 17.54 dBm (giving about 0.238 rad modulation index).
  • These numbers actually do not match expectations. I'm sure the modulation depth is correct as I cross-checked with sideband powers. So most probably the RF amplification is somehow getting attenuated.
  • However, the RF path looks fine when I did a TF analysis using Test Ports and used liso model to convert into how photodiode signal would behave. But this test didn't really go to high enough simulated photocurrent (to create photocurrent in the paths equivalent to 3 mW light, 144 V input is required).
  • Not sure why this is happening as the MAX4107ES chip on the RFPD is also new. All rails and voltage regulators are at nice values. There is no clipping as well as the PDH error signal is beautifully ideal.
  • But the PMC is locked robustly, so I'm gonna look past this for now.
  • 4.292 mW light is reaching the cavity.

South Path

  • South path's PMC was in good running condition even after the incident. So I didn't open the RFPD at all.
  • I did the same thing in this path, increased EOM driving signal to 17.54 dBm and measured PDH error signal.
  • This path has exceptionally attenuated RF response given this RFPD (SN001) was measured to have 14430 V/A AC transimpedance.
  • With 6.32 mW light falling on the RFPD, PDH error signal is 780 mVp-p.
  • The exact reason for low error signal is still a mystery, but the PMC is locking robustly with good looking PDH error signal.
  • The overlap at South PMC is poorer than North PMC because the incoming light has a lot of higher order modes.
  • 2.768 mW light is falling on the cavity. Clearly, this is not just higher more losses. More alignment is required on South PMC overlap so that more power can reach afterward.

Near future steps:

  • Do same analysis for FSS loops. Already have done some part in this regard.
  • Align the beams better everywhere.
  • Get beatnote and test beatnote detector.
  • Fix the actual ground loop noise issue!
ELOG V3.1.3-